### v1

# ProxNF: Neural Field Proximal Training for High-Resolution 4D Dynamic Image Reconstruction

Abstract
Accurate spatiotemporal image reconstruction methods are needed for a wide range of biomedical research areas but face challenges due to data incompleteness and computational burden. Data incompleteness arises from the undersampling often required to increase frame rates and reduce acquisition times, while computational burden emerges due to the memory footprint of high-resolution images with three spatial dimensions and extended time horizons. Neural fields, an emerging class of neural networks that act as continuous representations of spatiotemporal objects, have previously been introduced to solve these dynamic imaging problems by reframing image reconstruction to a problem of estimating network parameters. Neural fields can address the twin challenges of data incompleteness and computational burden by exploiting underlying redundancies in these spatiotemporal objects. This work proposes ProxNF, a novel neural field training approach for spatiotemporal image reconstruction leveraging proximal splitting methods to separate computations involving the imaging operator from updates of the network parameter. Specifically, ProxNF evaluates the (subsampled) gradient of the data-fidelity term in the image domain and uses a fully supervised learning approach to update the neural field parameters. By reducing the memory footprint and the computational cost of evaluating the imaging operator, the proposed ProxNF approach allows for reconstructing large, high-resolution spatiotemporal images. This method is demonstrated in two numerical studies involving virtual dynamic contrast-enhanced photoacoustic computed tomography imaging of an anatomically realistic dynamic numerical mouse phantom and a two-compartment model of tumor perfusion.
Index Terms:
Dynamic imaging, Neural fields, Proximal method, Stochastic optimization, Spherical Radon transform, Computer-simulation study
I Introduction
Dynamic image reconstruction with high spatial and temporal resolution is important for understanding the evolution of many biological systems [1, 2, 3]. In many dynamic image reconstruction problems, measurements are undersampled and frame-by-frame reconstruction methods yield unsatisfactory resolution. In such cases, higher resolution can be achieved by utilizing a spatiotemporal reconstruction method in which the image is treated as a single spatiotemporal object [4]. However, large-scale spatiotemporal image reconstruction can quickly become computationally prohibitive and exceed memory requirements in three spatial dimensions. Therefore, high-resolution dynamic image reconstruction needs to exploit underlying redundancies in the image to reduce the number of parameters necessary for storing the image [5, 6].
On the one hand, optimization-based image reconstruction that employ semiseparable or low rank approximations has often been advocated to address these challenges. Previous work has demonstrated the use of such approaches to enable several dynamic imaging applications, including X-ray tomography [7], positron electron tomography [8], magnetic resonance imaging[9, 6], and more recently also photoacoustic computed tomography [10, 11, 12]. On the other hand, the use machine learning, and neural fields in particular, has been actively investigated and shows great promise for the solution of this class of image reconstruction problems. Neural fields (NFs) are a particular class of neural networks that provide a continuous representation of an object in space and time. Using a neural field reduces a dynamic image reconstruction problem into estimating network weights by solving a nonconvex optimization problem [13, 14, 15]. Unlike other deep learning methods, NFs are self-supervised. The neural field model is trained only using the measurement data and does not require ground truth images, or labels[16]. Additionally, NFs allow for memory-efficient representation of the object [17, 18] and provide implicit regularization based on the neural architecture[19].
Previous works with NFs utilized first-order (accelerated) gradient based methods [18, 19, 20, 21, 22, 17, 23], such as Adam [24], to train the network parameters. However, training a NF for image reconstruction in this fashion requires the imaging operator and its adjoint to be computed at each gradient update, greatly slowing down training [13, 25, 14]. Furthermore, while computational efficient implementation of discrete-to-discrete imaging operators are readily available, the implementation of a continuous-to-discrete imaging operator acting on a NF representation incur severe computational and memory burden due to the need of repeated evaluation of the NF [13].
This work proposes ProxNF, a proximal splitting method for self-supervised training of a NF to enable memory-efficient, high-resolution spatiotemporal image reconstruction. The key contribution of this splitting method is a separation of concerns between the updates of NF parameters and calls to the imaging operator. In this approach, an update direction in the image domain is computed using the gradient of the data-fidelity term, which requires an evaluation of the discrete-to-discrete imaging operator and its adjoint. Once this update direction is formed, no further computations involving the imaging operator are needed. The NF parameters are then updated independently from the imaging operator by solving a supervised subproblem (the proximal operator). In essence, this method reframes the original self-supervised problem to solving a series of supervised subproblems while reducing the number of calls to the imaging operator and thus the computational burden. Furthermore, the update direction can be computed in a memory efficient manner with framewise-batched gradients of the data-fidelity term and imaging operator. ProxNF allows for reconstruction of large, high-resolution spatiotemporal images that were previously intractable. This method is demonstrated with dynamic image reconstructions in two and three spatial dimensions inspired by dynamic contrast-enhanced photoacoustic tomography imaging of tumor perfusion.
The remainder of this paper is structured as follows. In Section II, spatiotemporal image reconstruction is formulated as an optimization problem and reconstruction methods utilizing NFs for representing the dynamic object are introduced. In Section III, the proposed ProxNF method for training a NF for spatiotemporal image reconstruction is derived. This section also discusses a specific NF architecture, the partition of unity network, and introduces an adaptation to this architecture for representing high-resolution, high-dimensional dynamic images. In Section IV, the design of two numerical studies is presented along with the computational framework for virtual dynamic contrast-enhanced photoacoustic imaging of a numerical tumor perfusion phantom in two and three spatial dimension. The first study assesses the representation power of the proposed NF architecture applied to the dynamic tumor perfusion numerical phantom compared to a class of semiseparable approximations and to the original partition of unity network architecture. The second study compares dynamic image reconstruction in two and three spatial dimensions using the proposed NF representation with ProxNF and a model-based reconstruction method using nuclear norm regularization. In Section V, the results of these numerical studies are presented. Section VI presents the conclusions drawn from these results and future extensions.
II Background
II-A Spatiotemporal image reconstruction
II-A1 Continuous-to-discrete formulation
Image reconstruction aims to estimate an object $f$ based on a set of indirect measurements $\boldsymbol{d}$. The object $f$ belongs to an infinite dimensional Hilbert space $\mathbb{U}$. The measurements $\boldsymbol{d}$ belong to a finite dimensional space $\mathcal{D}$ and are associated to $f$ via a possibly nonlinear imaging operator $\mathcal{H}:\mathbb{U}\rightarrow\mathcal{D}$, also known as a forward operator, via
$\boldsymbol{d}=\mathcal{H}(f)+\boldsymbol{n},$ (1)
where $\boldsymbol{n}$ represents additive noise.
This works is concerned with the specific case of dynamic image reconstruction. That is, $f$ is a spatiotemporal function belonging to $\mathbb{U}=\mathbb{L}^{2}(\Omega_{T})$ with $\Omega_{T}=\Omega\times[0,T]$, where $\Omega\subset\mathbb{R}^{d}$ ($d=2$ or 3) denotes the spatial domain and $T$ is the time horizon.
In dynamic image reconstruction, the imaging operator $\mathcal{H}$ is the concatenation of imaging operators $\{\mathcal{H}_{k}\}_{k=1}^{K}$, one for each imaging frame $k=1,\ldots K$, where $K$ denotes the number of frames. Under the quasi-static approximation that the object $f$ does not change during the acquisition of the $k$-th imaging frame, the measurements $\boldsymbol{d}_{k}\in\mathcal{D}_{k}$ collected at the $k$-th frame are described by
$\boldsymbol{d}_{k}=\mathcal{H}_{k}(f(\cdot,t_{k}))+\boldsymbol{n}_{k},\quad k=1,\ldots,K,$ (2)
where $t_{k}=(k-1)\Delta T\in[0,T]$ is the sampling time of frame $k$, $\Delta T$ is the sampling interval, $f(\cdot,t_{k})$ is the object at frame $k$, and $\boldsymbol{n}_{k}$ represents measurement noise. For several applications of practical relevance the measurement aperture translates or rotates between imaging frames [4, 26, 27, 28, 12, 11], yielding imaging operators $\{\mathcal{H}_{k}\}_{k=1}^{K}$ that are different for each frame and measurements that are severely undersampled. In what follows, the notations
$\begin{array}[]{l}\displaystyle\boldsymbol{d}=\sum_{k=1}^{K}\boldsymbol{d}_{k}\otimes\mathbf{e}_{k},\quad\boldsymbol{n}=\sum_{k=1}^{K}\boldsymbol{n}_{k}\otimes\mathbf{e}_{k},\\ \displaystyle\mathcal{H}(f)=\sum_{k=1}^{K}\mathcal{H}_{k}(f(\cdot,t_{k}))\otimes\mathbf{e}_{k}\end{array}$ (3)
are introduced to write the dynamic imaging model in Eqn. (2) in the general form of Eqn. (1). Above, $\{\mathbf{e}_{k}\}_{k=1}^{K}$ indicates the canonical basis of $\mathbb{R}^{K}$ and the symbol $\otimes$ represents the outer product of two vectors yielding a rank-1 matrix.
Using the above notation, spatiotemporal image reconstruction can be formulated as an optimization problem
$\min_{f}\mathcal{J}(f):=\mathcal{L}(f)+\mathcal{R}(f).$ (4)
Above, the cost functional $\mathcal{J}:\mathbb{U}\rightarrow\mathbb{R}$ is the sum of the data-fidelity functional $\mathcal{L}:\mathbb{U}\rightarrow\mathbb{R}$, which is minimized when $\mathcal{H}(f)=\boldsymbol{d}$, and the regularization functional $\mathcal{R}:\mathbb{U}\rightarrow\mathbb{R}$, which encodes prior information on the sought-after image. For the case of Gaussian additive noise, the data-fidelity functional $\mathcal{L}(f)=\sum_{k=1}^{K}\mathcal{L}_{k}(f(\cdot,t_{k}))$ is given by
$\begin{array}[]{cl}\displaystyle\mathcal{L}(f)&=\frac{1}{2\sigma^{2}}\|\mathcal{H}(f)-\boldsymbol{d}\|^{2}\\ &=\sum_{k=1}^{K}\frac{1}{2\sigma^{2}}\|\mathcal{H}_{k}(f(\cdot,t_{k}))-\boldsymbol{d}_{k}\|^{2},\end{array}$
where the weight $\sigma^{2}$ represents the marginal variance of the additive noise $\boldsymbol{n}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})$.
II-A2 Discrete-to-discrete formulation
To derive the discrete-to-discrete formulation, the spatial domain $\Omega$ is discretized using a Cartesian grid with $M$ voxels. The coordinates of the voxel centers are denoted with $\{\boldsymbol{r}_{m}\}_{m=1}^{M}$. The spatial and temporal expansion functions $\beta_{m}(\boldsymbol{r})$ and $\gamma_{k}(r)$ are chosen as the indicator functions of the $m$-th spatial voxel ($m=1,\ldots,M$) and the $\left(t_{k}-\frac{1}{2}\Delta T,t_{k}+\frac{1}{2}\Delta T\right)$ time interval ($k=1,\ldots,K$), respectively. The finite dimensional subspace $\mathbb{U}_{M,K}\subset\mathbb{U}$ consists of all functions $f_{M,K}$ of the form
$f_{M,K}(\boldsymbol{r},t)=\sum_{k=1}^{K}\sum_{m=1}^{M}{\rm f}_{m,k}\beta_{m}(\boldsymbol{r})\gamma_{k}(t),$ (5)
where ${\rm f}_{m,k}\in\mathbb{R}$ ($m=1,\ldots,M$, $k=1,\ldots,K$) are the expansion coefficients. Eqn. (5) defines an isomorphism $\mathcal{I}:\mathbb{R}^{M\times K}\mapsto\mathbb{U}_{M,K}$ mapping the matrix $\mathbf{F}$ with entries $\{{\rm f}_{m,k}\}_{m,k}$ to the function $f_{M,K}(\boldsymbol{r},t)$. The inner product in $\mathbb{U}_{M,K}$ is defined as
[…]
The adjoint operator $\Pi_{M,K}^{*}:\mathbb{R}^{M,K}\mapsto\mathbb{U}_{M,K}\subset\mathbb{U}$ is defined with respect to the inner product in Eqn. (6) as
$\Pi_{M,K}^{*}(\mathbf{F})=\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{1}{V_{m,k}}{\rm f}_{m,k}\beta_{m}(\boldsymbol{r})\gamma_{k}(t).$ (7)
The notations $\mathbf{f}_{k}$ and $\Pi_{M,K}^{k}(f)$ are also introduced to denote $k$-th column of $\mathbf{F}$ and $\Pi_{M,K}(f)$, respectively.
The discrete-to-discrete imaging operators $\{\mathbf{H}_{k}:\mathbb{R}^{M}\rightarrow\mathcal{D}_{k}\}_{k=1}^{K}$ are constructed by evaluating their continuous-to-discrete counterpart acting on the spatial expansion functions $\{\beta_{m}(\boldsymbol{r})\}_{m}$. Specifically, they are defined as
$\mathbf{H}_{k}=\sum_{m=1}^{M}\mathcal{H}_{k}(\beta_{m}(\cdot))\otimes\mathbf{i}_{m},$ (8)
where $\{\mathbf{i}_{m}\}_{m=1}^{M}$ denotes the canonical basis of $\mathbb{R}^{M}$. The discrete-to-discrete spatiotemporal imaging operator $\mathbf{H}:\mathbb{R}^{M\times K}\mapsto\mathcal{D}$ is defined in a similar manner to its continuous-to-discrete counterpart, c.f. Eqn. (3). Its action on a matrix $\mathbf{F}$ representing the discretized spatiotemporal object function $f$ reads
$\mathbf{H}(\mathbf{F})=\sum_{k=1}^{K}\left(\mathbf{H}_{k}\mathbf{f}_{k}\right)\otimes\mathbf{e}_{k}.$ (9)
Using the above notation, the discrete counterpart of the optimization-based formulation of spatiotemporal image reconstruction reads
$\min_{\mathbf{F}\in\mathbb{R}^{M\times K}}\mathbf{J}(\mathbf{F})\coloneqq\mathbf{L}(\mathbf{F})+\mathbf{R}(\mathbf{F}),$ (10)
where
$\mathbf{L}(\mathbf{F})=\sum_{k=1}^{K}\mathbf{L}_{k}(\mathbf{f}_{k})=\frac{1}{2\sigma^{2}}\sum_{k=1}^{K}\|\mathbf{H}_{k}\mathbf{f}_{k}-\boldsymbol{d}_{k}\|^{2}$
and
$\mathbf{R}(\mathbf{F})=\mathcal{R}(\mathcal{I}(\mathbf{F})).$
II-B Neural fields
A neural field (NF) is a neural network that provides a representation of continuous objects, such as images, videos, and audio signals [20]. An NF can be described as a neural network $\Phi_{\boldsymbol{\xi}}:\Omega_{T}\rightarrow\mathbb{R}$ that acts as a map between spatiotemporal coordinates to the values of an image at those coordinates. NF architectures represent a finite-dimensional, nonlinear topological subspace of $L^{2}(\Omega_{T})$ parametrized by a set of network parameters $\boldsymbol{\xi}\in\mathbb{R}^{N_{p}}$. For many classes of images, in particular biomedical images[23, 29], NF can provide accurate representations using a fixed network architecture and requiring a relatively small number of parameters $N_{p}$ compared to the number of voxels needed to represent the image with a comparable resolution [17]. NFs are then able to significantly reduce the memory requires for storing high-resolution, high-dimensional images.
Given a target function $f\in\mathbb{L}^{2}(\Omega_{T})$, an NF can be trained to represent $f$ by solving the supervised embedding problem
$\min_{\boldsymbol{\xi}}\frac{1}{2}\|f-\Phi_{\boldsymbol{\xi}}\|_{2}^{2},$ (11)
where $\|f-\Phi_{\boldsymbol{\xi}}\|_{2}$ denotes the $L^{2}(\Omega_{T})$ norm of $f-\Phi_{\boldsymbol{\xi}}$. Once trained, the NF can be evaluated at arbitrary points in $\Omega_{T}$ and, thus, rendered at any chosen resolution.
II-B1 Spatiotemporal image reconstruction using neural fields
In the case of spatiotemporal image reconstruction, the object $f$ is not directly accessible. Previous works have demonstrated that NFs can be directly trained using measurement data alone by replacing the sought-after object $f$ with its NF representation $\Phi_{\boldsymbol{\xi}}$ into the optimization problem in Eq. (4)[13, 14]. This gives rise to the self-supervised training problem
$\min_{\boldsymbol{\xi}}\mathcal{J}(\Phi_{\boldsymbol{\xi}})\coloneqq\frac{1}{2\sigma^{2}}\|\mathcal{H}(\Phi_{\boldsymbol{\xi}})-\boldsymbol{d}\|^{2}+\mathcal{R}(\Phi_{\boldsymbol{\xi}}).$ (12)
Advantages of using an NF for spatiotemporal image reconstruction in Eq. (12) include reduced memory requirements for optimization. However, this does come with a few disadvantages. First, optimizing over the space of network parameters to train the NF is a nonconvex optimization problem, even in the case when the original problem in Eq. (4) is convex. The nonconvex nature of this problem means there are no guarantees on finding a global minimizer; additionally non-convex problems are in general more difficult to solve than convex problems. Second, while computationally efficient implementations of several canonical discrete-to-discrete imaging operators are readily available, the evaluation of a continuous-to-discrete imaging operator acting on the NF representation, $\mathcal{H}(\Phi_{\boldsymbol{\xi}})$ is computationally expensive and further slows down training.
II-B2 Partition of unity networks
One particular NF network architecture is the partition of unity network (POUnet) [30], which has previously been utilized for spatiotemporal image reconstruction in two spatial dimensions [13]. The POUnet consists of three components: a trainable partition of unity $\boldsymbol{\Psi}_{\boldsymbol{\eta}}:\Omega_{T}\rightarrow[0,1]^{P}$ parameterized by a set of weights $\boldsymbol{\eta}\in\mathbb{R}^{W}$, a trainable coefficient matrix $\boldsymbol{C}\in\mathbb{R}^{L\times P}$, and a fixed set of spatiotemporal basis functions $\boldsymbol{B}:\Omega_{T}\rightarrow\mathbb{R}^{L}$. Specifically, the POUnet NF representation reads
$\Phi_{(\boldsymbol{\eta},\boldsymbol{C})}(\mathbf{x})=\boldsymbol{\Psi}_{\boldsymbol{\eta}}^{T}(\mathbf{x})\boldsymbol{C}^{T}\boldsymbol{B}(\mathbf{x}),$
where the partition of unity network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}(\mathbf{x})=(\psi_{\boldsymbol{\eta}}^{1}(\mathbf{x}),\ldots,\psi_{\boldsymbol{\eta}}^{P}(\mathbf{x}))$ is such that
$\sum_{p=1}^{P}\psi^{p}_{\boldsymbol{\eta}}(\mathbf{x})=1,\text{ and }\psi^{p}_{\boldsymbol{\eta}}(\mathbf{x})\geq 0.$
By letting $\boldsymbol{\xi}=(\boldsymbol{\eta},\boldsymbol{C})$, an approximate solution to the minimization problem in Eq. (12) can be found using an alternating minimization strategy
$\begin{array}[]{c}\boldsymbol{C}^{i+1}=\operatorname*{argmin}_{\boldsymbol{C}}\mathcal{J}(\Phi_{(\boldsymbol{\eta}^{i},\boldsymbol{C})})\\ \boldsymbol{\eta}^{i+1}=\operatorname*{argmin}_{\boldsymbol{\eta}}\mathcal{J}(\Phi_{(\boldsymbol{\eta},\boldsymbol{C}^{i+1})}).\end{array}$ (13)
The POUnet architecture is advantageous because of this fast optimization strategy, particularly in the case of convex cost functionals $\mathcal{J}$ in which updates of the coefficients $\boldsymbol{C}$ are also convex. Furthermore, the POUnet architecture can provide theoretical accuracy guarantees based on a priori error estimates for piecewise polynomial interpolation [30].
III Method
This section presents the main contribution of this work: the development of a proximal training method for NF-based spatiotemporal image reconstruction, ProxNF. In the proposed method, all computations involving the imaging operator are separated from updates of the NF parameters. Furthermore, ProxNF leverages preexisting discrete implementations of the imaging operator and fast NF embedding methods to scale to high-resolution, high-dimensional imaging problems. This work also develops a modified partition of unity NF architecture that is capable of representing dynamic images with three spatial dimensions.
III-A ProxNF: a proximal training method for neural field dynamic image reconstruction
Previously, it has been shown that NFs can be utilized for small-scale spatiotemporal image reconstruction problems and trained with gradient based methods[13, 14]. However, these gradient-based methods become very slow when scaling to high-resolution, high-dimensional problems. This lack of scalability is due to the difficulties outlined in Section II-B1 of nonconvexity and expensive imaging operator evaluations.
Notably, these disadvantages are greatly reduced in the supervised embedding problem in Eq. (11) for a few reasons. First although this problem is nonconvex, embedding is very well studied and great amounts of effort have been made to design gradient-based training methods that can approximately find a “good” local minimum. Here a “good” local minimum refers to a local minimum with a sufficiently low suboptimality gap. Second, evaluating the embedding training objective and its gradient require no calls to the imaging operator and is computationally inexpensive. This means that the supervised problem can be solved relatively quickly and easily when compared to Eq. (12).
To leverage the existing capabilities for solving NF embedding problems, this work proposes a novel formulation of the spatiotemporal image reconstruction problem as constrained minimization problem, in which the sought-after object estimate $f^{*}$ is constrained to the manifold $\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}\coloneqq\{f\in L^{2}(\Omega_{T})|f={\Phi}_{\xi},\ \xi\in\mathbb{R}^{W}\}$ of all functions that can be represented exactly by the chosen NF architecture. That is, the object estimate $f^{*}$ is defined as
$f^{*}\coloneqq\operatorname*{argmin}_{f\in\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}\frac{1}{2\sigma^{2}}\left\|\mathcal{H}(f)-\boldsymbol{d}\right\|^{2}+\mathcal{R}(f).$ (14)
In this work, a stochastic generalized plug-and-play method, referred to as ProxNF, is developed to approximately solve Eqn. (14), as outlined in Algorithm 1. Plug-and-play methods [31, 32, 33, 34] are a generalization of the proximal gradient algorithm [35, 36] in which the proximal operator is replaced by some denoising method. To this aim, a generalized proximal operator (denoiser) that computes a regularized projection of a spacetime object estimate $f\in L^{2}(\Omega_{T})$ onto the manifold $\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}$ is introduced. Its evaluation requires the solution of a regularized embedding problem defined as
$\operatorname{Prox}_{\alpha,\mathcal{R}}^{\textnormal{NF}}(f):=\operatorname*{argmin}_{\tilde{f}\in\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}\frac{1}{2\alpha}\|\tilde{f}-f\|^{2}+\mathcal{R}(\tilde{f}).$ (15)
An update direction $g\in\mathbb{U}$ can then be computed, for example using gradient of the data-fidelity term, in the image domain $\mathbb{U}$ and independent of the NF architecture. With the proximal operator, the updated estimate can then be projected back to the NF manifold and regularized. This can be summarized in the iterative process
$g^{n}=\nabla_{f}\mathcal{L}(f^{n}),\ f^{n+1}=\operatorname{Prox}_{\alpha^{n},\mathcal{R}}^{\textnormal{NF}}(f^{n}-\alpha^{n}g^{n}),$ (16)
where $f^{n}$ represents the image estimate at the current iteration, $g^{n}$ the corresponding update direction at the $n$-th iteration, and $\alpha^{n}>0$ is the (possibly) iteration dependent step-length.
This method then reframes the original self-supervised problem as a series of, easier to solve, supervised subproblems. In this work the update direction $g$ is chosen to be the gradient of the data-fidelity term, but could be generalized to include other choice of update direction (potentially conjugate gradient or (Quasi-)Newton directions).
Approximating the continuous-to-discrete the imaging operator $\mathcal{H}(f)$ with $\boldsymbol{H}(\Pi_{M,K}f)$, the update direction $g$ in Eq. (16) is numerically computed as
$g^{n}=\Pi_{M,K}^{*}\nabla_{\mathbf{F}}\mathbf{L}(\Pi_{M,K}f^{n}),$
where $\Pi_{M,K}^{*}:\mathbb{R}^{M\times K}\rightarrow\mathbb{L}^{2}(\Omega_{T})$ is the adjoint of $\Pi_{M,K}$ defined in Eqn (7). However, explicitly forming this update direction $g^{n}$ requires to store an expansion coefficient matrix $\mathbf{G}^{n}\in\mathbb{R}^{M\times K}$ and thus can compromise the memory efficiency provided by the NF representation of the object. Memory efficiency can be preserved by computing an update direction $\tilde{g}$ using stochastically subsampled frames. For a stochastically selected collection of frames $\{k_{1},\ldots,k_{J}\}$ (with $J\ll K$), the update direction is then computed as
$\tilde{g}^{n}=\Pi_{M,K}^{*}\sum_{j=1}^{J}\nabla_{\mathbf{f}_{k_{j}}}\mathbf{L}_{k_{j}}(\Pi_{M,K}^{k_{j}}f^{n})\otimes\mathbf{e}_{k_{j}},$
which only requires storing an expansion coefficient matrix $\tilde{\mathbf{G}}^{n}\in\mathbb{R}^{M\times J}$. This proposed method is outlined in Algorithm 1.
It is worth noting that Nesterov acceleration is not used Algorithm 1. This is because proximal splitting methods combining randomized evaluation of the data-fidelity gradient with Nesterov acceleration lack theoretical convergence guarantees and may be susceptible to instabilities due to accumulation of errors in the momentum term [37, 38, 39, 40].
Initialize neural field $f^{0}=\Phi_{\boldsymbol{\xi}}$.
for $n=0,1,2,\ldots$ do
       Randomly select collection of frames $\{k_{1},\ldots,k_{J}\}.$ for $j=1,\ldots,J$ do
             Compute frame-wise coefficients
$\mathbf{f}_{k_{j}}^{n}=\Pi_{M,K}^{k_{j}}f^{n}$
Compute frame-wise directions
$\mathbf{g}_{k_{j}}^{n}=\nabla_{\mathbf{f}_{k_{j}}}\mathbf{L}_{k_{j}}(\mathbf{f}_{k_{j}}^{n}).$
       end for
      Form the update direction
$\tilde{g}^{n}=\Pi_{M,K}^{*}\sum_{j=1}^{J}\mathbf{g}_{k_{j}}^{n}\otimes\mathbf{e}_{k_{j}}.$
For a step size $\alpha$, apply proximal update
$f^{n+1}\leftarrow\operatorname{Prox}^{\textnormal{NF}}_{\alpha,\mathcal{R}}(f^{n}-\alpha\tilde{g}^{n})$
end for
Algorithm 1 ProxNF: Dynamic image reconstruction with neural fields via proximal splitting
III-B Adapted partition of unity neural fields
This work proposes utilizing modified partition of unity network of the form
$\Phi_{(\boldsymbol{\eta},\boldsymbol{C})})(\boldsymbol{r},t)=\boldsymbol{\Psi}^{T}_{\boldsymbol{\eta}}(t)^{T}\boldsymbol{C}^{T}\boldsymbol{B}(\boldsymbol{r}),$
where the partition of unity network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ only depends on the time coordinate $t$ and the basis functions $\boldsymbol{B}$ only depends on the spatial coordinate $\boldsymbol{r}$. In particular, $\boldsymbol{B}(\boldsymbol{r})=\begin{pmatrix}\beta_{1}(\boldsymbol{r})&\ldots&\beta_{M}(\boldsymbol{r})\end{pmatrix}$ shares the same basis functions $\{\beta_{m}\}_{m=1}^{M}$ in Eqn. (5). This representation allows for the high spatial resolution provided by traditional discretization methods and smooth temporal transitions while still exploiting spatiotemporal redundancies. Moreover, this architecture greatly simplifies discrete projection operations. That is for a selected frame $k$,
$\Pi_{M,K}^{k}\Phi_{(\boldsymbol{\eta},\boldsymbol{C})}=\mathbf{C}\boldsymbol{\Psi}_{\boldsymbol{\eta}}(t_{k}).$
In this work, the regularization term $\mathcal{R}$ was chosen to be Tikhonov regularization on the on the spatial derivatives computed with finite difference and on the temporal derivatives computed with automatic differentiation of the learned partition of unity. This choice of regularization then allowed the proximal $\operatorname{Prox}_{\alpha,\mathcal{R}}^{\textnormal{NF}}$ to be computed using the alternating minimization algorithm outlined in Eq. (13), where the coefficient updates can be accomplished by solving a linear system while updates of the partition update can be updated using a stochastic gradient method.
IV Numerical Studies
This work utilized two numerical studies to assess the proposed NF representation and proposed ProxNF method for spatiotemporal image reconstruction. These studies were inspired by dynamic contrast-enhanced (DCE) photoacoustic computed tomography (PACT) imaging of tumor perfusion in small animal models [41]. PACT is a hybrid biomedical imaging modality that combines benefits from both optical and ultrasound imaging[42]. Specifically, PACT combines the benefits of high soft tissue contrast from optical imaging and high resolution detection principles from ultrasound imaging to create high resolution maps of optical properties. PACT image formation utilizes a short laser pulse in the near infrared range to illuminate a biological tissue. The biological tissue then absorbs the optical energy and generates heat and a local increase in pressure, which is described as the photoacoustic effect. This pressure distribution then propagates as an acoustic wave where it can be measured outside the tissue using ultrasound detectors. Under the assumption of a non-lossy homogeneous acoustic medium, that is, a medium with homogeneous speed of sound and density and no acoustic attenuation, the relationship between the induced pressure distribution and the data can be modeled using the circular Radon transform in two spatial dimensions and the spherical Radon transform in three spatial dimensions [43, 44]. The numerical studies presented in this work focus on estimation of this spatially and temporally varying induced pressure distribution from sparse-view measurements collected from a circular aperture in two spatial dimensions and a cylindrical aperture in three spatial dimensions.
IV-A Construction of the dynamic object
The dynamic objects used in these studies were spatiotemporal maps of the spatially and temporally varying induced pressure distribution within a numerical tumor perfusion model in two and three spatial dimensions. The anatomical structures of these objects were constructed utilizing an abdominal section extracted from an anatomically realistic whole-body murine numerical phantom (MOBY) [45, 46]. MOBY comprises detailed three-dimensional (3D) anatomical structures featuring tens of organs and models physiologically realistic cardiac and respiratory motions. A spherical lesion of diameter 7 mm was inserted in the abdominal region at a depth of 5 mm to mimic a subcutaneous tumor flank. To simulate DCE imaging of a tumor perfusion model, the concentration of an optical contrast agent in the lesion and blood-filled organs was computed using a two compartment Toft’s model [47, 48, 49]. In this model, the contrast agent was introduced after 90 seconds and injected over a 30 second window following a slow bolus pattern. Optical properties (optical absorption coefficient and scattering coefficient) corresponding to a 800 nm laser pulse wavelength were then assigned to each tissue type based on physiologically plausible chromophore concentrations (oxygenated hemoglobin, deoxygenated hemoglobin, water, fat, contrast agent)[50]. Contrast agent optical properties were chosen to mimic those of indocyanine green (ICG) [51], in which case the absorption spectrum has a peak near the chosen wavelength.
The diffusion approximation[52] was used to simulate photon transport at each imaging frame. The incident fluence was assumed constant on the surface of the object. The fluence distribution within the object was obtained by solving the partial differential equation stemming from the diffusion approximation using the open-source finite element library FEniCS [53], thus yielding a volumetric map of dynamic induced pressure distribution at each imaging frame. For the experiments in two spatial dimensions a slice was extracted from the 3D induced pressure phantom.
Figure 1: Numerical dynamic mouse abdominal phantom extract from mouse whole body phantom (MOBY) and shown at beginning and and of inspiration during a 5 second breathing period. The phantom is comprised of detailed three-dimensional (3D) anatomical structures featuring several of organs and models physiologically realistic cardiac and respiratory motions. A spherical lesion of diameter 7 mm was inserted in the abdominal region at a depth of 5 mm to mimic a subcutaneous tumor flank and displayed in dark red.
IV-B Definition of the virtual imaging systems
IV-B1 Dynamic imaging system for two spatial dimensions
Figure 2 illustrates the virtual imaging system used in the numerical study with two spatial dimensions. The measurements geometry was circle of radius $R$ (shown in green), along which two groups of 5 idealized point-like ultrasound transducers are distributed. The sensors in each group were spaced $1^{\circ}$ apart on the transducer arc. Each group of sensors was distributed antipodal from one another. The object was placed at the center of the imaging system and fully contained within a square region of size $L$ (field of view). At each imaging frame, data were collected simultaneously by each sensor corresponding to the integral along $I$ concentric arcs. Sensors were rotated by an angle $\Delta\Theta$ after each imaging frame. The dynamic changes of the object were assumed to be slow enough such that the object was not expected to vary significantly during the acquisition of each imaging frame (quasi-static approximation). This assumption is often valid for several dynamic PACT applications since the duration of the light excitation pulse at each imaging frame is in the range of a few nanoseconds while the dynamic changes in the object have a time-scale in the range of milliseconds to seconds.
Under these assumptions, the imaging operator in two spatial dimensions can then be described by the circular Radon transform (CRT). For a sensor location $r_{0}^{(k)}\in\mathbb{R}^{2}$, measurement radius $\ell$, and imaging frame $t_{k}$, the CRT can be expressed as the integral along a circular arc
[…]
The superscript $k$ in $r_{0}^{(k)}$ is used to highlight the fact that the location $r_{0}^{(k)}$ of the sensor changes at each imaging frame. The discrete-to-discrete counterpart of the CRT imaging operator was implemented using the AirToolsII MatLab toolbox [54] and stored as a compressed sparse column matrix[55]. The imaging systems parameters are displayed in Table I.
Figure 2: Illustration of the virtual system for dynamic imaging using a CRT model. At each imaging frame, data are collected by idealized point-like transducers (shown in blue). Transducers rotate of an angle $\Delta\Theta$ between frames.
IV-B2 Dynamic imaging system for three spatial dimensions
For the numerical study in three spatial dimensions, the measurement geometry was a cylinder of radius $R$ and height $A$, along which four groups of idealized point-like ultrasound transducers were distributed. Each group of sensors consisted of 5 columns of sensors spaced $1^{\circ}$ apart. Each column consisted of 96 sensors spaced $\Delta z$ apart. These four groups of sensors were evenly spaced $90^{\circ}$ apart.
The object was placed at the center of the imaging system and fully contained within a rectangular prism region of side length $L$ in the $x$ and $y$ direction and height $H$. At each imaging frame, data were collected at each sensor ($I$ measurements per sensor). Sensors rotated an angle $\Delta\Theta$ along the z-axis of the cylindrical geometry at each frame. Again the dynamic changes in the object were assumed to be slow enough such that the object is approximately constant at each imaging frame.
Under these assumptions, the imaging operator in three spatial dimensions can be described by the spherical Radon transform (SRT). The SRT, which computes the integral of the object function over concentric spherical shells, is the 3D generalization of the CRT. To construct the discrete-to-discrete counterpart of the SRT operator, this work utilizes a computationally efficient method, described in [56], which exploits the symmetry of the cylindrical measurement geometry and decomposes the SRT as the composition of two 2D CRT operators [57]. The imaging system parameters for these experiments are reported in Table I.
TABLE I: Dynamic imaging system parameters
Imaging system
FOV diameter $L$ 3.72 cm
FOV height $H$ (3D) 2 cm
Aperture $R$ 2.63 cm
Aperture height $A$ 4 cm
Acquisition time $T$ 648 s
Number of frames $K$ 1,296
Sensor separation $1^{\circ}$
Rotation angle $\Delta\Theta$ $5^{\circ}$
Number of sensors per group $5$
Two spatial dimensions
Number of pixels $M=M_{s}^{2}$ $186^{2}$
Rings per view $I$ 263
Relative noise $0.04$
Number of sensor groups 2
Number of views per frame $S$ $10$
Total number of sensors 10
Three spatial dimensions
Number of voxels $M=M_{s}^{2}M_{z}$ $186^{2}\times 100$
Rings per view $I$ 323
Relative noise $0.01$
Number of sensor groups 4
Number of views per frame $S$ $20$
Number of sensors per column $Z$ 96
Total number of sensors 1,920
IV-C Study Design
Study 1: Dynamic embedding
The first numerical study solved the embedding problem and assessed the proposed NF architecture for representing the dynamic object described in Section IV-A. For the case involving two spatial dimensions, the proposed NF architecture was compared to the original NF architecture with polynomial basis functions [13] and a low-rank (semiseparable) decomposition of the dynamic object. The proposed POUnet architecture used a network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ with 4 hidden layers of width 140 nodes and one output layer of width $P=10$ partitions. Sinusoidal and softmax activation functions were employed in the hidden and output layers, respectively. The set of basis function $\mathbf{B}$ consisted of $L=186^{2}=34596$ pixel indicator functions. This resulted in a representation with $427,020$ parameters. The original POUnet architecture used a network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ with 5 hidden layers of width 280 nodes and output layer of width $P=100$ partitions. Third order space-time separable polynomials were employed as basis functions ($L=40$). This resulted in a representation with $425,800$ parameters. These two NF representations were compared to a semiseparable approximation of rank $r=12$, SS(12), which has $430,704$ parameters. Comparatively, storing the dynamic object as a collection of spatiotemporal voxels requires $\sim 20$ million parameters. Therefore these representations allow for a 40-fold reduction in the number of necessary parameters.
For the case involving three spatial dimensions, the proposed NF architecture was compared to a low-rank, semiseparable, decomposition of the dynamic image. The POUnet architecture again utilized the same network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ as the 2D case and a set of basis function $\mathbf{B}$ that contained $L=186^{2}\cdot 100=3,459,600$ voxel indicator functions. This resulted in a representation with $34,988,424$ parameters. The semiseparable approximation used a rank $r=10$, SS(10), which has $34,608,960$ parameters. These representations yield over two-order of magnitude reduction in the number of parameters compared to representing the object as a collection of spatiotemporal voxels values ($\sim 5$ billion parameters).
Study 2: Dynamic reconstruction
The second numerical study explored dynamic image reconstruction in two (CRT imaging operator) and three (SRT imaging operator) spatial dimensions. In both cases, two spatiotemporal image reconstruction approaches were utilized. The first reconstruction technique utilized the proposed NF architecture and the proposed ProxNF method in Algorithm 1. The second reconstruction technique, which serve as a reference, was spatiotemporal image reconstruction with nuclear norm regularization (STIR-NN) [58]. In particular, for the case involving three space dimensions, the memory-efficient maximal-rank-constrained method in [11] was implemented. Measurement data were corrupted with i.i.d. additive Gaussian noise with zero mean and standard deviation $\sigma$ controlled by a specified relative noise level (RNL), defined as $\operatorname{RNL}=\sigma/\|\boldsymbol{d}\|_{\infty}$. In the two spatial dimensions case, the RNL was set to 0.04 and, in the three spatial dimensions case, the RNL was set to 0.01. Regularization weights for all reconstructions were chosen using Morozov’s discrepancy principle [59] in which the regularization weight was swept so that the standard deviation of the data-fidelity offset was approximately equal to the noise standard deviation. In the two spatial dimensions case CRT data were collected using 2 rotating sets of sensors (5 sensors per set) placed antipodally apart. In the three spatial dimensions case, SRT data were generated from 4 rotating sets of sensors placed 90 degree apart from each other.
Evaluation metrics
In these studies, image quality was quantified using relative root mean square error (RRMSE) and structural similarity index measure (SSIM) [60], evaluated both on the entire image and on a region of interest containing the lesion. Image quality was also assessed by analyzing the estimated time activity curve, which is the average signal over a volume of interest (the lesion).
V Results
The numerical studies in two spatial dimensions were performed on a workstation with two Intel Xeon Gold 5218 Processors (16 cores, 32 threads, 2.3 GHz, 22 MB cache each), 384 GB of DDR4 2933Mhz memory, and one NVidia Titan RTX 24GB graphic processing unit (GPU). The numerical studies in three spatial dimensions were performed on the Delta supercomputer at the National Center for Supercomputing Applications (NCSA) of the University of Illinois at Urbana-Champaign. Specifically, a single node with dual AMD Milan processors (64 core, 128 threads, 2.55 GHz, 96MB cache each), 256 GB of DDR4 3200 MHZ memory and two quad A100 40GB GPUs was used in the studies below.
The proposed method, ProxNF, was implemented using PyTorch, an open-source machine learning framework [61], and executed on GPUs. The POUnet parameters $\boldsymbol{\eta}$ and $\boldsymbol{C}$ were initialized by performing an embedding of the time-averaged object for the representation studies and by solving a static image reconstruction problem on time-averaged measurements for the reconstruction studies.
Following initialization, the NFs in the representation study were trained using the alternating minimization approach in Eq. (13). Similarly after initialization, the NFs in the reconstruction study were trained with the proposed proximal splitting method in Algorithm 1. A reduction of the order of magnitude in the norm of the proximal gradient was used as stopping criterion. The batch size was $J=324$ and time frames were stochastically selected at each iteration. Tikhonov regularization was used to penalize the $L^{2}$ norm of the gradient of the spatiotemporal image. The regularized embedding problem at each proximal update was solved as follow. The linear systems corresponding to the updates of $\boldsymbol{C}$ were solved using the conjugate gradient method. The internal updates for $\boldsymbol{\eta}$ were accomplished using an Adam optimizer [24]. Specifically, 10,000 iterations with a fixed learning rate of $10^{-5}$ and batch size of 100,000 randomly sampled spatiotemporal points were used.
The spatiotemporal reconstructions in two spatial dimensions using the reference method (STIR-NN) were accomplished utilizing the Fast Iterative Shrinkage Algorithm (FISTA) [62] as implemented in UNLocBoX, a MATLAB toolbox for proximal-splitting methods [63]. The low rank semiseparable approximation of the image with two spatial dimensions was accomplished using SVD command in MATLAB and soft-thresholding. The spatiotemporal reconstructions in three spatial dimensions using the reference method (STIR-NN) were accomplished using the memory-efficient randomized algorithm in [11] with a maximum rank constraint of 32. The randomized singular value decomposition [64] was employed to compute the low rank decomposition and enforce the maximum rank constraint.
V-A Study 1: Dynamic embedding problem
In this study, the NF representations and semiseparable approximations of rank $r$, SS($r$), were compared for a 2D cross-sectional slice and 3D volume extracted from the dynamic tumor perfusion phantom described in Section IV-A.
Figure 3: Study 1: Average lesion activity of true object in two spatial dimensions (blue), SS(12) approximation (green), original POUnet architecture (purple), and proposed POUnet architecture (red). The proposed POUnet architecture better captures the dynamic changes of the object compared to the original POUnet architecture in [13].
Figure 4: Study 1: Dynamic object in two spatial dimensions (top), Rank 12 semiseparable approximation SS($12$) (second row), neural field using original POUnet architecture (third row), and neural field using proposed POUnet architecture are shown at selected frames throughout time period. The dynamic range of the grayscale image is $041e-$ 04 to $0.06$. A logarithmic scale was used in the color map to increase visualization contrast at depth. An animation of each representation constructed in this study is available in the supplemental multimedia materials (Video 1).
Figure 5: Study 1: Average lesion activity of true object in three spatial dimensions (blue), the semi-separable rank 10 approximation (green), and the neural field representation (red). The semiseparable and neural field representation lead to an almost identical, smooth lesion activity curve that underestimates activity at later times.
Figure 6: Study 1: Dynamic object in three spatial dimensions (top), Rank 10 semiseparable approximation SS($10$) (second row), neural field representation (bottom row) with proposed POUnet architecture are shown at selected frames throughout time period. Paraview [65] was used for volume rendering. Dynamic range of the rendering is [0-0.06] using a linear scale. Both methods provide high fidelity representations but do slightly underestimate the tumor activity in later time periods. An animation of each representation constructed in this study is available in the supplemental multimedia materials (Video 2).
TABLE II: Study 1: Dynamic embedding problem. RRMSE, SSIM, ROI RRMSE, ROI SSIM, and LAC RRMSE achieved by the NF representations with original and proposed POUnet architectures and the rank $r$ semiseparable approximations SS($r$) in two (2D) and three (3D) spatial dimensions.
# of RRMSE SSIM ROI ROI LAC
Params RRMSE SSIM RRMSE
SS(12) 2D 430 K $0.0426$ $0.9964$ $0.0842$ $0.9114$ $0.0282$
Original NF 2D 425 K $0.2600$ $0.9399$ $0.2571$ $0.8432$ $0.0591$
Proposed NF 2D 427 K $0.0670$ $0.9943$ $0.1005$ $0.9026$ $0.0291$
SS(10) 3D 35 M $0.0558$ $0.9983$ $0.1661$ $0.9864$ $0.0988$
Proposed NF 3D 35 M $0.0571$ $0.9982$ $0.1661$ $0.9862$ $0.1174$
Two spatial dimensions
The accuracy metrics—RRMSE and SSIM over the entire image, RRMSE and SSIM over the region of interest containing the inserted lesion, and the lesion activity curve RRMSE— achieved by each representation are reported in Table II. The lesion activity curves associated with these representations and with the object are shown in Fig. 3. Selected frames of these representations and of the object are shown in Fig. 4. The proposed NF architecture showed superior approximation properties (as quantified by the above accuracy metrics) compared to the original POUnet architecture in [13], and comparable accuracy compared to the semiseparable approximation with a similar number of parameters.
Three spatial dimensions
The accuracy metrics for the proposed POUnet representation and the semiseparable approximation of rank 10 (SS(10)) are reported in Table II. The average lesion activity curves associated with these representations and with the object are shown in Fig. 5. Selected frames of these representations and the object are shown in Fig. 6. The proposed NF architecture achieved comparable accuracy metrics to those achieved by the semiseparable approximation with a similar number of parameters.
V-B Study 2: Dynamic reconstruction
In this study, the proposed ProxNF method was assessed and compared against a well-established spatiotemporal image reconstruction method with nuclear norm regularization (STIR-NN). Specifically, a 2D cross-sectional slice and 3D volume extracted from the dynamic tumor perfusion phantom employed in Study 1 were virtually imaged using the stylized photoacoustic tomography system described in Section IV-B.
TABLE III: Study 2: Dynamic reconstruction. Number of parameters, RRMSE, SSIM, ROI RRMSE, ROI SSIM, and LAC RRMSE achieved by the proposed ProxNF method and the reference STIR-NN method in two (2D) and three (3D) spatial dimensions.
# of RRMSE SSIM ROI ROI LAC
Params RRMSE SSIM RRMSE
STIR-NN 2D 5 M $0.2467$ $0.8822$ $0.2530$ $0.7826$ $0.1406$
ProxNF 2D 427 K $0.2392$ $0.9039$ $0.2404$ $0.8279$ $0.0791$
STIR-NN 3D 52 M $0.2915$ $0.9493$ $0.4021$ $0.9085$ $0.3193$
ProxNF 3D 35 M $0.2452$ $0.9665$ $0.2288$ $0.9564$ $0.1916$
Two spatial dimensions
The number of parameters required to represent the reconstructed estimate of the object and the corresponding accuracy metrics—RRMSE and SSIM for the entire image, the RRMSE and SSIM for the region of interest with the inserted lesion, and the lesion activity curve (LAC) RRMSE—are reported in Table III. As displayed by RRMSE and SSIM over the entire image and the region of interest, the ProxNF reconstruction was more accurate compared to the STIR-NN method. Additionally, the ProxNF reconstruction led to a more accurate estimation of the lesion activity, as indicated by the LAC RRMSE and illustrate in Fig. 7. Furthermore, the ProxNF reconstruction required less than a tenth of the parameters used by the STIR-NN. Selected frames of these reconstructions and the original object are shown in Fig. 8.
Three spatial dimensions
The number of parameters and accuracy metrics for the ProxNF and STIR-NN methods are reported in Table III. The reconstructed time activities for the inserted lesion are shown in Fig. 9. Selected frames of these reconstructions and the original object are shown in Fig. 10. Note that the ProxNF reconstructions is more accurate in terms of RRMSE and SSIM, both over the entire image and the region of interests, and provides a more accurate time activity curve.
Figure 7: Study 2: Average lesion activity of the dynamic object in two spatial dimensions (blue), nuclear norm reconstruction (green), and neural field reconstruction (red). Note how ProxNF is more accurate in the pre- and post-injection phases.
Figure 8: Study 2: Spatiotemporal image reconstruction in two spatial dimensions. Dynamic object (top) and its estimates produced by STIR-NN (middle), and ProxNF (bottom) are shown at selected imaging frames. The dynamic range of the grayscale image is $[10^{-04},0.06]$. A logarithmic scale was used in the color map to increase visualization contrast at depth. An animation of each reconstructed estimate is available in the supplemental multimedia materials (Video 3). ProxNF can produce an object estimate with comparable accuracy to that produced by STIR-NN, while reducing the memory requirements of one order of magnitude.
Figure 9: Study 2: Average lesion activity of true object in three-spatial dimensions(blue) and neural field reconstruction (red). The neural field representation underestimates the true lesion activity but matches its behavior.
Figure 10: Study 2: Spatiotemporal image reconstruction in three spatial dimensions. Paraview [65] was used for volume rendering. Dynamic object (top) and its estimates produced by STIR-NN (middle), and ProxNF (bottom) are shown at selected imaging frames. Dynamic range of the rendering is [0-0.05] using a linear scale. An animation of each reconstruction from this study is available in the supplemental multimedia materials (Video 4). The proposed method can better capture the dynamic changes in the object compared to the reference method (STIR-NN). Furthermore, as shown in Video 4, ProxNF can resolve the respiratory motion of the numerical phantom while the reconstructed estimate produced reference method only captures a time-averaged dynamic over the respiratory cycle.
VI Conclusion
This works proposes ProxNF, a proximal splitting method for high-resolution, high-dimensional image reconstruction using a neural field (NF) representation. NFs are a special class of neural networks that provide accurate representations of continuous objects and can be utilized for memory-efficient image reconstruction. NFs have demonstrated benefits in dynamic image reconstruction, including automatic exploitation of spatiotemporal redundancies[20, 17, 13, 14]. The proposed ProxNF method increases the scalability of NF-based image reconstruction approaches enabling high-resolution, long time horizon volumetric dynamic imaging. ProxNF achieves this by separating computations involving updates of the NF parameters from those involving the imaging operator, thereby reducing the number of evaluations of the imaging operator and allowing for efficient training of NF parameters. The efficacy of the proposed partition-of-unity-based NF architecture and ProxNF method were illustrated in two numerical studies motivated by dynamic contrast-enhanced photoacoustic computed tomography imaging applications.
The first study showed that the proposed partition of unity network (POUnet) architecture is able to accurately represent a dynamic images in two and three spatial dimensions, achieving an accuracy comparable to that yielded by a space-time semiseparable approximation using a similar number of parameters. The study also showed that the proposed architecture has better representation properties than the original POUnet architecture in [13].
The second study considered dynamic image reconstruction problems in two and three spatial dimensions. A numerical tumor perfusion phantom was virtually imaged with a limited number of views per imaging frame. This study demonstrated that the NF architecture trained using ProxNF outperforms a well-established image reconstruction method utilizing nuclear norm regularization in terms of RRMSE, SSIM, while using a smaller number of parameters to represent the sought-after image. Furthermore, the NF method yielded a more accurate estimate of the lesion activity curve compared to that achieved by the method using nuclear norm regularization.
In summary, this work established the feasibility of NF representations for dynamic image reconstruction of high-resolution, high-dimensional images. ProxNF, the proposed proximal splitting method, helps overcome the computational shortcomings of previous approaches to spatiotemporal image reconstruction using NFs [13]. A key advantage of the proposed method is that it is agnostic of the chosen neural representation. The plug-and-play formulation of ProxNF allows to seamlessly replace the POUnet architecture used in this work with any other state-of-the-art neural field architecture[19, 17, 18, 20, 21, 13]. The primary limitation of this work is the use of numerical phantoms, future studies should investigate the application of the proposed method to in-vivo (clinical or preclinical) data. Additionally, another limitation is the relatively long computational time, however, training the NF can be accelerated by use of modern machine learning software[66]. Future work may extend the proposed method to 5D (three spatial dimensions, one time dimension, and one wavelength dimensions) quantitative PACT, in which dynamic changes in the object optical properties are estimated from multispectral PACT data by use of a non-linear multiphysics imaging operator coupling photon transport and wave propagation.
Appendix A Code and Data Availability
The python code implementing the proposed dynamic image reconstruction method using neural field is available from [67] under GPLv3. The absorption coefficient and induced pressure maps of the dynamic object used in the numerical studies are available from [68] under CC-0 public dedication.
Appendix B Supplementary Multimedia Material
Supplementary multimedia materials include the following. Video 1 depicts the results of the representation study in two spatial dimensions displayed in Fig. 4 alongside the time activity curves for the inserted lesion. Video 2 depicts the representation results in three spatial dimensions displayed in Fig. 6 alongside the time activity curves and the underlying anatomy. Video 3 depicts the results of the dynamic image reconstruction study in two spatial dimensions displayed in Fig. 8 alongside the time activity curves for the inserted lesion. Video 4 depicts the dynamic image reconstruction results in three spatial dimensions displayed in Fig. 10 alongside the time activity curves and the underlying anatomy.